OpenStudy (anonymous):

Find the y if the slope is 1/2 in (-1, 1) (5,y)

OpenStudy (mertsj):

Don't you want to try this one now? You solve it and so will I and we can see if we agree.

OpenStudy (anonymous):

y-1=6 1/2?

OpenStudy (anonymous):

Oh wait

OpenStudy (mertsj):

\[\frac{1}{2}=\frac{y-1}{5+1}\]

OpenStudy (anonymous):

y-3=6

OpenStudy (mertsj):

\[\frac{1}{2}=\frac{y-1}{6}\]

OpenStudy (anonymous):

*y-1=6

OpenStudy (anonymous):

y=4?

OpenStudy (mertsj):

Yes. y = 4. Way to go!!!

OpenStudy (mertsj):

Isn't learning fun???

OpenStudy (anonymous):

Oh very. Sarcasm intended haha

OpenStudy (mertsj):

Well, just be thankful that God gave you a good brain so that you can learn. Pity those unfortunate people who can't.

OpenStudy (anonymous):

What if in this case, you find x and its -3-7/4-x and the slope is -1?

OpenStudy (anonymous):

Yes, praise the Lord. :P

OpenStudy (mertsj):

i would cross multiply. I think that might be easiest.

OpenStudy (anonymous):

x=6?

OpenStudy (mertsj):

\[\frac{-10}{4-x}=\frac{-1}{1}\]

OpenStudy (mertsj):

I got -6

OpenStudy (anonymous):

I'll try it again

OpenStudy (mertsj):

-3-7=-10

OpenStudy (mertsj):

And slope is -1 which is -1 over 1 or 1 over -1

OpenStudy (mertsj):

When you cross multiply you get -4+x=-10

OpenStudy (anonymous):

Okay so the -1 has to multiplied like this?: -1(-4-x)

OpenStudy (mertsj):

exactly.

OpenStudy (mertsj):

No. it's -1(4-x)

OpenStudy (anonymous):

Oops my bad :P

OpenStudy (anonymous):

okay I get it thank you!

OpenStudy (mertsj):

yw